OpenStudy (anonymous):

EMERGENCY PLEASE HELP WILL FAN AND MEDAL The regular price of a child's entry ticket to a water park is $9 less than that for an adult's. The park offers half off all entry tickets during the off-peak season. The Sandlers paid a total of $132 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket: 132 = one-halfx + 2(x − 9) What is the regular price of a child's ticket?

OpenStudy (anonymous):

ok so in total, it was 139 right?

OpenStudy (anonymous):

WHAT WAS

OpenStudy (anonymous):

the total price they payed

OpenStudy (anonymous):

NO IT WAS 132

OpenStudy (anonymous):

child = c adult = a total = t c = a - 9 t = 132 t = 1/2a + 2c t = 1/2a + 2(a - 9) 1/2a + 2a - 18 = 132 5a = 228 a = $45.60 Therefore child = 45.60 - 9.00 = $36.60

OpenStudy (anonymous):

hold on. im sorry, i gave you the wrong answer.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

That equation is correct. 132 = 1/2x + 2(x - 9) Solve for x to get the regular price of an adult ticket. 132 = 1/2x + 2x - 18 150 = 2.5 x 60 = x minus 9 is 51, the answer is 51

OpenStudy (anonymous):

THANKS SO MUCH

OpenStudy (anonymous):

no problem.